1.In class of 40 students, 38 offer maths, 24 offer English. Each of the students offer at least one of the two subjects. How many students offer both subjects?

2. If A={ 3 ,4 ,5 ,9 and B={1,2,6,9,7}. Find (I) A–B (ii)B n A.
1
Expert's answer
2019-11-04T10:02:31-0500

1.In the class 40 students, 24 offer English and each of students offer at least one of the two subjects. It means that "40-24=16" students offer Maths. But we know that 38 students offer Maths. 16 students are not enough 22 to 38 ("38-16=22" ). But all other student (expect 16 who offer Maths) offer English. It means 22 students offer 2 subjects.

Answer: 22 students offer both subjects.

2. "A= \{ 3,4,5,9 \},B= \{ 1,2,6,9,7 \}" .

(I) "A-B" means that from the set A you need to pick up the common elements between A and B or "A-B= \{ x \in A \mid x \notin B \}" . The common element between A and B is 9. So "A-B = \{ 3,4,5 \}".

(II) "B \cap A = \{ x \in A \mid x \in B \}" or the common elements A and B. It's 9. "B \cap A = \{ 9\}" .

Answer: "A-B = \{ 3,4,5 \}" , "B \cap A = \{ 9\}" .